OpenStudy (anonymous):

Produce a 97% confidence interval for ρ if the proportion in a sample of size 320 is 0.58.

OpenStudy (anonymous):

see table in the attachment

OpenStudy (kropot72):

The standard error of the sample proportion is \[\sqrt{\frac{p(1-p)}{n}}=you\ can\ calculate\] When you have found the value of the standard error, the 97% confidence interval is: \[0.58\pm(2.17\times\ standard\ error)\]